SpaceRef

SpaceRef


SATOPS Morning Report - March 13, 2003† Day 072†

Status Report From: China National Space Administration
Posted: Thursday, March 13, 2003

DMSP†

DMSP operations were nominal over the past 24 hours.†

POES†

The telemetry and command databases for all polar operational spacecraft were distributed and installed on March 12 to the SOCC, Fairbanks, and Wallops.† The next database distribution and installation to all sites is scheduled for Wednesday April 16.† An apparent problem with the distribution to Fairbanks is being reviewed today.†

NOAA-12 rev 61435 / F at 1308Z on March 12:† Call up pass to monitor telemetry and review effects resulting from the MSU scan motor commanding, which was accomplished at 0953Z on March 12.† All systems (including spacecraft power balance) appear to be nominal.†

NOAA-15 rev 25097 / F at 1932Z on March 12:† During this pass, GAC2B played down from the spacecraft without incident, but the data could not be extracted from the IMUX during post-pass.† Fairbanks suspected that the problem was caused by a bad board in the IMUX recorder.† This data recovery was unsuccessfully attempted on rev 25098 / F at 2111Z, due in part by a conflict with a DMSP operational support.† At 2235Z, 100% of GAC2B was recovered at WCDAS.†

GOES†

Correction to yesterday's report:† In addition to the GOES-8 image missed on March 11 due to comm controller switching activities, GOES-10 missed one image and one sounding.†

At 0245Z on March 13, Central inquired about the 0245 full disk.† The RPM archive only contained fragments of the image.† The crew noted that the image activity started and ended on-time on the spacecraft.†

Eclipse operations for all spacecraft continue.† Eclipse schedules for GOES-8 and GOES-10 (which can be used to track the day-to-day image deletions due to KOZ and eclipse) can be found at:† http://www.oso.noaa.gov/goes/koz-eclipse/koz-eclipse.htm

GOES-12 remains in a test phase, in normal on-orbit mode, drifting at an eastward rate of ~0.33 degrees per day toward the GOES-East position where it will replace GOES-8 in April.† Routine SXI imaging is proceeding along with SXI test exposures.†

GOES-9 is currently at approximately 173.6 W drifting at a westward rate of ~0.735 degrees per day toward 155 E longitude where it will support the GMS-5 replacement mission in April.† Canberra and Goldstone DSN ranging activities are continuing.

POES operations were nominal over the past 24 hours.

// end //

More status reports and news releases or top stories.

Please follow SpaceRef on Twitter and Like us on Facebook.

SpaceRef Newsletter